# Photovoltaic power forecasting using quantum machine learning

### v1

@@ Abstract
Predicting solar panel power output is crucial for advancing the energy transition but is complicated by the variable and non-linear nature of solar energy. This is influenced by numerous meteorological factors, geographical positioning, and photovoltaic cell properties, posing significant challenges to forecasting accuracy and grid stability. Our study introduces a suite of solutions centered around hybrid quantum neural networks designed to tackle these complexities. The first proposed model, the Hybrid Quantum Long Short-Term Memory, surpasses all tested models by over $40\%$ lower mean absolute and mean squared errors. The second proposed model, Hybrid Quantum Sequence-to-Sequence neural network, once trained, predicts photovoltaic power with $16\%$ lower mean absolute error for arbitrary time intervals without the need for prior meteorological data, highlighting its versatility. Moreover, our hybrid models perform better even when trained on limited datasets, underlining their potential utility in data-scarce scenarios. These findings represent a stride towards resolving time series prediction challenges in energy power forecasting through hybrid quantum models, showcasing the transformative potential of quantum machine learning in catalyzing the renewable energy transition.
@@ I Introduction
Electricity generation prediction, especially for photovoltaic (PV) systems, is a crucial tool for renewable energy adoption sabri2022novel ; li2020review . The global economy must radically reduce emissions to stay within the $1.5\degree$C pathway (Paris Agreement) and the transition to renewable energy sources is necessary to achieve these objectives ellabban2014renewable . According to the IEA, solar PV’s installed power capacity is poised to surpass that of coal by 2027, becoming the largest in the world.
@@
In this article, we present three types of hybrid quantum models as potential solutions for PV power forecasting. We assess the performance of our proposed models using a publicly accessible dataset, encompassing a comprehensive array of meteorological variables as well as hourly mean PV power measurements spanning a $21$-month period. This dataset, along with the data preprocessing and analytical methodologies employed, is described in detail in Section II.1.
Our first proposed HQNN architecture, articulated in Section II.2, incorporates classical fully connected layers with a vanilla variational repetitive quantum layer (VVRQ). Our second model, delineated in Section II.3, constitutes a hybrid quantum adaptation of the classical recurrent neural network, termed the Hybrid Quantum Long Short-Term Memory with quantum depth-infused layer (HQLSTM). While the first two models can predict the power for a certain hour ahead, the third model, presented in Section II.4, a Hybrid Quantum Sequence-to-Sequence Neural Network with quantum depth-infused layer, HQSeq2Seq, after training is capable of forecasting PV power for arbitrary time intervals without requiring prior meteorological data.
Remarkably, despite having fewer parameters, our hybrid quantum models outperform their classical counterparts in terms of more accurate predictions, including trained on a reduced dataset. We summarize our conclusions and outline future research directions in Section III.
@@ II Results
The application of HQNNs in addressing time series prediction challenges, specifically in forecasting PV power output offers several advantages. Primarily, their capability to operate within an exponentially larger computational search space enables them to efficiently capture intricate data patterns and relationships kordzanganeh2023exponentially . This feature not only enhances forecast accuracy biamonte2017quantum ; senokosov2023quantum but also streamlines the learning process, requiring fewer iterations for model optimization asel1 . Furthermore, the inherent capacity of quantum technologies to manage the uncertainty and noise ubiquitous in data offers more resilient and trustworthy predictions schuld2018implementing . This is particularly pertinent to power forecasting, given the inherent noise in meteorological data. Recent research also suggests that quantum models can be represented as partial Fourier series, positioning them as potential universal function approximators schuld2021effect , thereby broadening their applicability and efficacy in predictive tasks.
In terms of architecture, an HQNN is an amalgamation of classical and quantum components. The classical segments may consist of fully connected layers, convolutional layers, or recurrent layers, while the quantum segments are typically represented by variational quantum circuits (VQCs) or their contemporary modifications kordzanganeh2023parallel ; sedykh2023quantum .
@@ II.1 Dataset
To underscore the advantages of hybrid quantum models using empirical evidence, we selected a publicly accessible dataset malvoni2016data from a conventional generation plant situated in the Mediterranean region. This dataset not only provides comprehensive data but also allows benchmarking with results from various algorithms available in literature. A comparative analysis of our model’s predictions and those from the study by kaloop2021novel is provided in Section II.5. The dataset, presented as a numerical table showcased in Fig. 1, encompasses variables like hourly mean ambient temperature ($T_{a}$), hourly mean module temperature ($T_{m}$), hourly mean solar irradiance recorded on two tilted planes with tilt angles of 3 and 15 degrees ($I_{3},\,I_{15}$), and hourly mean PV power ($P$) spanning $21$ months, accounting for more than $500$ days.
@@
To ensure the validity and precision of the data, meticulous preprocessing and analysis were undertaken. We discovered approximately 20 anomalies in the original dataset. To maintain a continuous timeline, missing data points were replaced with the arithmetic mean of the preceding and succeeding day’s values. Additionally, data corresponding to the date “12/31/13” was excluded as it contained all-zero values, suggesting an error in data collection. As a result, we obtained an uninterrupted dataset ranging from 4:00 AM on “3/5/12” to 12:00 AM on “12/30/13”.
Additional in-depth analysis of the dataset was also conducted for a more nuanced understanding. Fig. 2(a) delineates the hourly distribution of PV power across the entire period. As expected, peak PV power values occur during midday, whereas night time values plummet to zero. Fig. 2(b) portrays monthly PV power fluctuations, which are more volatile compared to daily patterns, likely attributable to the limited number of full-year periods in the dataset. Fig. 2(c) presents a correlation matrix for the dataset features, identifying solar irradiances $I_{3}$ and $I_{15}$ as the features most correlated with PV power. Finally, the joint distribution of dataset features depicted in Fig. 2(c) further confirms that solar intensity is the feature most highly correlated with PV power.
Figure 3: The architectures of: (a) Hybrid Quantum Neural Network with VVRQ layer, (b) Hybrid Quantum Long Short-Term memory, (c) QDI layer used in HQLSTM model, (d) Hybrid Quantum Seq2Seq with QDI layer.
@@ II.2 HQNN
This section introduces our first proposed model, referred to as the HQNN. As illustrated in Fig. 1, the model accepts weather data spanning $24$ consecutive hours as its input. The output is a prediction of the PV power for the upcoming $25$th hour. The HQNN presented at the Fig. 3(a) is a combination of classical fully-connected layers, in our case with $120$, $17$ and $8$ neurons, and a VVRQ layer, which is a VQC, consisting of $q$ qubits and $d$ repetitions of variational layers, each distinguished by unique weights. The choice of $120$ neurons is methodical: the model ingests $5$ distinct features for each of the $24$ hours, resulting in a total of $120=5*24$. The determination of the remaining parameters stemmed from an extensive hyperparameter optimization process, detailed in the subsequent sections.
Initially, every qubit in the VVRQ layer is set to the state $\left|0\right\rangle$. We subsequently encode the classical data by converting it into rotation angles around one of the $X$, $Y$, $Z$ axes using $R_{x}$, $R_{y}$, $R_{z}$ gates respectively. This conversion employs the angle embedding technique PhysRev.70.460 . For each qubit, the rotation angle, denoted by $x_{j}$, is determined by the $j$-th component of the input vector.
Following this, the variational layer is applied, which can either utilize “basic” or “strong” entanglements. For the “basic” entanglement, each qubit undergoes a rotation by an angle $w_{j}^{i}$ around the $X$ axis, subsequently followed by a layer of CNOT gates Barenco_1995 . Conversely, for the “strong” entanglement, each qubit is sequentially rotated by the angles $w^{(Z_{1})}_{ji}$, $w^{(Y_{2})}_{ji}$, and $w^{(Z_{3})}_{ji}$ around the $Z$, $Y$, and $Z$ axes, respectively. This sequence is then followed by a layer of CNOT gates. In both cases, the variables $i$ and $j$ play crucial roles in determining the operations. The variable $i$ signifies the particular wire to which the operation is applied, and it takes values from the set ${1,2,\dots,q}$. Meanwhile, the variable $j$ represents the number of variational layers and ranges from ${1,2,\dots,d}$.
Lastly, all qubits are measured in Pauli-$Z$ basis, yielding the classical vector $v\in\mathbb{R}^{q}$. This output serves as input for a subsequent classical fully-connected layer. This layer processes information from $q$ neurons into $1$ neuron that predicts the power value.
@@ II.3 HQLSTM
This section presents a description of our second hybrid model – HQLSTM, which is a hybrid analog of the classical LSTM model hochreiter1997long , with which predictions will be compared in the following sections. LSTM architectures have garnered significant attention in the realm of time series forecasting, including in predicting PV power gao2023short ; sharadga2020time ; tovar2020pv .
HQLSTM models have proven themselves well for solving problems from various fields. Examples of successful use of this model are the tasks of natural language processing wang2022adverse , the detection of software vulnerabilities akter2022automated , and predicting solar radiation yu2023prediction .
In this proposed model we added a quantum layer to each of the LSTM gates chen2020quantum . Let’s take a closer look at our implementation, depicted in Fig 3(b). The input to the model is:
@@
Subsequently, this concatenated vector is partitioned into four distinct groups, for the four gates of the LSTM cell. As illustrated in Fig. 3(b-c), each group is directed to the input of its corresponding quantum layer, symbolized by the QDI square.
The outputs from QDI layers are transformed via classical fully-connected layers to standardize their dimensions to $h_{dim}$. Following this, activation functions together with appropriate for each of the $4$ gates transformations, similar to the classical LSTM, are applied to the outputs originating from the quantum layers. This processing culminates in the derivation of the new cell state $C(t)$ and the hidden state $h(t)$ vectors.
The process operates in a cyclical manner. For each iteration, the vector from the current time step, $x(t)$, and the hidden vector from the previous step, $h(t-1)$, serve as inputs to the HQLSTM. This iterative process is executed as many times as the input width; in our case input width equals $24$. Subsequently, all the hidden vectors are concatenated to produce a single composite vector. This vector is then processed through a fully-connected layer consisting of a single neuron, which outputs a value that predicts the PV power.
In our first proposed architecture, the HQNN, the quantum layer functioned as a vanilla layer, where variational layers were sequentially placed after the encoding layer. In contrast, in the HQLSTM approach we used a QDI layer pharma as depicted in Fig. 3(b). Here, variational layers are positioned multiple times before the encoding layer (green rectangular) and additionally (purple rectangular), within each encoding layer (blue rectangular) to increase the layer’s expressivity.
@@ II.4 HQSeq2Seq
Here we present a hybrid version of the Sequence-to-Sequence (Seq2Seq) model, first introduced in sutskever2014sequence . Seq2seq models are widely used in natural language processing tasks kalchbrenner2013recurrent , where the length of input and output sequence is not pre-determined and can be variable. We can also apply the principle of Seq2Seq models to the power prediction task mu2023improved . That means we can feed the neural network with time series with arbitrary length and prompt it to give us the forecast for any hours ahead. In this problem setting, the longer the input time series is, the better the model prediction is. The same applies to the required output length: the shorter it is, the easier it is for the model to generate the forecast.
The seq2Seq model is a type of encoder-decoder model. The encoder is given the entire input sequence, which it uses to generate a context vector. This vector is used as an input hidden state for the decoder, so it literally provides it with “context”, according to which the decoder will generate the forecast. Thereby, the hidden dimensions of the encoder and the decoder must match. The decoder creates the output sequence step by step. It starts with only one entry: the one which is the last known. Based on this entry and the context vector, the decoder generates the second entry and appends it to the existing one. Now, the obtained two-entry sequence is once again fed into the decoder to generate the third entry. Then, the cycle repeats until the length of the generated sequence matches the length requested by the user.
@@
In the QDI layer, instead of attempting to use a qubit for each feature asel1 , we employed the data re-uploading technique schuld2021effect ; data_reuploading . Specifically, we work with $4$ qubits and structure them into a lattice of depth $4$ (depicted as a blue big rectangular in Fig. 3(d)). Each of our $16$ input features leading to the quantum layer is intricately encoded within this lattice. The first four features are mapped onto the initial depth, followed by the subsequent features in blocks of four. Encoding these classical features into the quantum domain, we adopt the “angle embedding” using $R_{z}$ gate. This operation effectively translates the input vector into a quantum state that symbolizes the preceding classical layer’s data.
@@
Entangling variational layers, signified by purple squares, are interposed between every encoding layer, ensuring optimal Fourier accessibility. Each variational layer has two components: rotations governed by trainable parameters, and sequential CNOT gates. The rotations are implemented by quantum gates that metamorphose the encoded input in line with the variational parameters, while the CNOT operations handle the entanglement of the qubits, facilitating quantum superposition.
@@
Each lattice depth, represented by each blue square, encompasses a variational layer (purple square). Moreover, prior to all encoding layers, we introduce a variational layer (designated by a green square) for enhanced model representation. Consequently, the total weight count in the quantum segment of our network is $20$. In the measurement phase, except for the first qubit, all qubits execute a CNOT operation targeting the first qubit, ensuring the $Y$-measurement spans all qubits. Therefore, the quantum layer’s output serves as the power value prediction for a specific hour.
@@
The input size of the encoder and decoder can differ, which is a substantial benefit. For instance, we can use all of the $5$ features to create a context vector, but request to generate the forecast for only $1$ feature. Exploiting this advantage, we will feed the Seq2Seq model with a window of all known features and demand the forecast only for the “Power” one.
For simplicity’s sake, we will train both models with fixed input and output length of $96$ hours and then try to vary the length in the testing stage.
@@
Figure 4: Results of training and testing of HQNN & MLP and HQLSTM & LSTM models. (a-b) Training and testing history of models. Solid and dotted lines represent training and testing of models respectively. Filled space shows standard deviation of models that were averaged over different testing subsets. Histogram for reduced training dataset. (c) Histogram shows mean MAE and RMSE metrics for our models on testing subset, averaged over different testing subsets. (f-g) Train and test learning curves for the Seq2Seq and HQSeq2Seq models. (h) Example of the classical and hybrid Seq2Seq models inference on the testing data. The models get $124$ hours of data as an input (before the dashed line) and gives forecast of “Power” into $137$ hours ahead (after the dashed line). The solid black line represents the ground truth value of the “Power” feature. MSE/MAE errors in this particular example are $0.0052/0.0349$ for Seq2Seq and $0.0040/0.0292$ for HQSeq2Seq.

### v2

@@ Abstract
Predicting solar panel power output is crucial for advancing the transition to renewable energy but is complicated by the variable and non-linear nature of solar energy. This is influenced by numerous meteorological factors, geographical positioning, and photovoltaic cell properties, posing significant challenges to forecasting accuracy and grid stability. Our study introduces a suite of solutions centered around hybrid quantum neural networks designed to tackle these complexities. The first proposed model, the Hybrid Quantum Long Short-Term Memory, surpasses all tested models by achieving mean absolute errors and mean squared errors that are more than $40\%$ lower. The second proposed model, the Hybrid Quantum Sequence-to-Sequence neural network, once trained, predicts photovoltaic power with $16\%$ lower mean absolute error for arbitrary time intervals without the need for prior meteorological data, highlighting its versatility. Moreover, our hybrid models perform better even when trained on limited datasets, underlining their potential utility in data-scarce scenarios. These findings represent progress towards resolving time series prediction challenges in energy forecasting through hybrid quantum models, showcasing the transformative potential of quantum machine learning in catalyzing the renewable energy transition.
quantum machine learning, hybrid quantum neural network, quantum depth infused layer, quantum LSTM, solar energy, photovoltaic power, time series
@@ I Introduction
Electricity generation prediction, especially for photovoltaic (PV) systems, is a crucial tool for renewable energy adoption sabri2022novel ; li2020review . The global economy must radically reduce emissions to stay within the $1.5\degree$C pathway (Paris Agreement) and the transition to renewable energy sources is necessary to achieve these objectives ellabban2014renewable . According to the International Energy Agency, solar PV’s installed power capacity is poised to surpass that of coal by 2027, becoming the largest in the world.
@@
In this article, we present three types of hybrid quantum models as potential solutions for PV power forecasting. We assess the performance of our proposed models using a publicly accessible dataset, encompassing a comprehensive array of meteorological variables as well as hourly mean PV power measurements spanning a $21$-month period. This dataset, along with the data preprocessing and analytical methods employed, is described in detail in Section II.1.
Our first proposed HQNN architecture, articulated in Section II.2, incorporates classical fully connected layers with a vanilla variational repetitive quantum layer (VVRQ). Our second model, delineated in Section II.3, constitutes a hybrid quantum adaptation of the classical recurrent neural network, termed the Hybrid Quantum Long Short-Term Memory with quantum depth-infused layer (HQLSTM). While the first two models can predict the power for a certain hour ahead, the third model, presented in Section II.4, a Hybrid Quantum Sequence-to-Sequence Neural Network with quantum depth-infused layer, HQSeq2Seq, after training, is capable of forecasting PV power for arbitrary time intervals without requiring prior meteorological data.
Remarkably, despite having fewer parameters, our hybrid quantum models outperform their classical counterparts in terms of more accurate predictions, including when trained on a reduced dataset. We summarize our conclusions and outline future research directions in Section III.
@@ II Results
The application of HQNNs in addressing time series prediction challenges, specifically in forecasting PV power output, offers several advantages. Primarily, their ability to operate within an exponentially larger computational search space enables them to efficiently capture intricate data patterns and relationships kordzanganeh2023exponentially . This feature not only enhances forecast accuracy biamonte2017quantum ; senokosov2024quantum but also streamlines the learning process, requiring fewer iterations for model optimization asel1 . Furthermore, the inherent capacity of quantum technologies to manage the uncertainty and noise ubiquitous in data offers more resilient and trustworthy predictions schuld2018implementing . This is particularly pertinent to power forecasting, given the inherent noise in meteorological data. Recent research also suggests that quantum models can be represented as partial Fourier series, positioning them as potential universal function approximators schuld2021effect , thereby broadening their applicability and efficacy in predictive tasks.
In terms of architecture, an HQNN is an amalgamation of classical and quantum components. The classical segments may consist of fully connected layers, convolutional layers, or recurrent layers, while the quantum segments are typically represented by variational quantum circuits (VQCs) or their contemporary modifications kordzanganeh2023parallel ; sedykh2024hybrid .
@@ II.1 Dataset
To investigate the advantages of hybrid quantum models, we selected a publicly accessible dataset malvoni2016data from a conventional generation plant situated in the Mediterranean region. This dataset not only provides comprehensive data but also enables benchmarking with results from various algorithms available in literature. A comparative analysis of our model’s predictions and those from the study by kaloop2021novel is provided in Section II.5. The dataset, presented as a numerical table showcased in Fig. 1, includes variables like: hourly mean ambient temperature, $T_{a}$; hourly mean module temperature, $T_{m}$; hourly mean solar irradiance recorded on two tilted planes with tilt angles of 3 and 15 degrees, $I_{3}$ and $I_{15}$; and hourly mean PV power, $P$, spanning $21$ months, accounting for more than $500$ days.
@@
To ensure the validity and precision of the data, careful preprocessing was performed. We discovered approximately 20 anomalies in the original dataset. To maintain a continuous timeline, missing data points were replaced with the arithmetic mean of the preceding and succeeding day’s values. Additionally, data corresponding to the date “12/31/13” was excluded since it contained all-zero values, suggesting an error in the data collection. As a result, we obtained an uninterrupted dataset ranging from 4:00 AM on “3/5/12” to 12:00 AM on “12/30/13.”
Additional in-depth analysis of the dataset was also conducted for a more nuanced understanding. Fig. 2(a) delineates the hourly distribution of PV power across the entire period. As expected, peak PV power values occur during midday, while night time values plummet to zero. Fig. 2(b) portrays monthly PV power fluctuations, which are more volatile compared to daily patterns, likely attributable to the limited number of full-year periods in the dataset. Fig. 2(c) presents a correlation matrix for the dataset features, identifying solar irradiances $I_{3}$ and $I_{15}$ as the features most correlated with PV power. Finally, the joint distribution of dataset features depicted in Fig. 2(c) further confirms that solar intensity is the feature most highly correlated with PV power.
Figure 3: The architectures of: (a) Hybrid Quantum Neural Network with a VVRQ layer, (b) Hybrid Quantum Long Short-Term memory, (c) QDI layer used in HQLSTM model, (d) Hybrid Quantum Seq2Seq with QDI layer.
@@ II.2 HQNN
This section introduces our first proposed model, referred to as the HQNN. As illustrated in Fig. 1, the model accepts weather data spanning $24$ consecutive hours as its input. The output is a prediction of the PV power for the upcoming $25$th hour. The HQNN presented in Fig. 3(a) is a combination of classical fully-connected layers, in our case with $120$, $17$ and $8$ neurons, and a VVRQ layer, which is a VQC, consisting of $q$ qubits and $d$ repetitions of variational layers, each distinguished by unique weights. The choice of $120$ neurons is methodical: the model ingests $5$ distinct features for each of the $24$ hours, resulting in a total of $120=5\times 24$. The determination of the remaining parameters stemmed from an extensive hyperparameter optimization process, detailed in the subsequent sections.
Initially, every qubit in the VVRQ layer is set to the state $\left|0\right\rangle$. We subsequently encode the classical data by converting it into rotation angles around one of the $X$, $Y$, $Z$ axes using $R_{x}$, $R_{y}$, $R_{z}$ gates, respectively. This conversion employs the angle embedding technique PhysRev.70.460 . For each qubit, the rotation angle, denoted by $x_{j}$, is determined by the $j$-th component of the input vector.
Following this, the variational layer is applied, which can either generate “basic” or “strong” entanglement. For the “basic” entanglement, each qubit undergoes a rotation by an angle $w_{j}^{i}$ around the $X$ axis, followed by a layer of CNOT gates Barenco_1995 . For the “strong” entanglement, each qubit is sequentially rotated by the angles $w^{(Z_{1})}_{ji}$, $w^{(Y_{2})}_{ji}$ and $w^{(Z_{3})}_{ji}$ around the $Z$, $Y$ and $Z$ axes, respectively. This sequence is then followed by a layer of CNOT gates. In both cases, the variables $i$ and $j$ play crucial roles in determining the operations. The variable $i$ signifies the particular wire to which the operation is applied, and it takes values from the set ${1,2,\dots,q}$. Meanwhile, the variable $j$ represents the number of variational layers and ranges from ${1,2,\dots,d}$.
Lastly, all the qubits are measured in Pauli-$Z$ basis, yielding the classical vector $v\in\mathbb{R}^{q}$. This output serves as input for a subsequent classical fully-connected layer. This layer processes information from $q$ neurons into $1$ neuron that predicts the power value.
@@ II.3 HQLSTM
This section presents a description of our second hybrid model – HQLSTM, which is a hybrid analog of the classical LSTM model hochreiter1997long , with which predictions will be compared in the following sections. LSTM architectures have garnered significant attention in the realm of time series forecasting, including in predicting PV power gao2023short ; sharadga2020time ; tovar2020pv . HQLSTM models have performed well on tasks such as natural language processing wang2022adverse , detecting software vulnerabilities akter2022automated , and predicting solar radiation yu2023prediction .
In this proposed model, we added a quantum layer to each of the LSTM gates chen2020quantum . Our implementation is depicted in Fig 3(b). The input to the model is:
@@
Subsequently, this concatenated vector is partitioned into four distinct groups for the four gates of the LSTM cell. As illustrated in Fig. 3(b-c), each group is directed to the input of its corresponding quantum layer, symbolized by the Quantum Depth-Infused (QDI) square.
The outputs from QDI layers are transformed via classical fully-connected layers to standardize their dimensions to $h_{dim}$. Following this, similar to the classical LSTM, activation functions together with transformations appropriate to each of the four gates, are applied to the outputs originating from the quantum layers. This processing produces the new cell state and hidden state vectors, $C(t)$ and $h(t)$, respectively.
The process operates in a cyclical manner. For each iteration, the vector from the current time step, $x(t)$, and the hidden vector from the previous step, $h(t-1)$, serve as inputs to the HQLSTM. This iterative process is executed as many times as the input width; in our case, the input width equals $24$. Subsequently, all the hidden vectors are concatenated to produce a single composite vector. This vector is then processed through a fully-connected layer consisting of a single neuron, which outputs a value that predicts the PV power.
In our first proposed architecture, the HQNN, the quantum layer functioned as a vanilla layer, where variational layers were sequentially placed after the encoding layer. In contrast, in the HQLSTM approach, we used a QDI layer pharma as depicted in Fig. 3(c).
In order to increase the expressivity, variational layers are positioned both before and after the encoding layer. The first variational layer is illustrated with green rectangles, the encoding layer with blue rectangles, and the final variational layer with purple rectangles.
@@ II.4 HQSeq2Seq
Here we present a hybrid version of the Sequence-to-Sequence (Seq2Seq) model, first introduced in sutskever2014sequence . Seq2seq models are widely used in natural language processing tasks kalchbrenner2013recurrent , where the length of input and output sequence is not predetermined and can be variable. We can also apply the principle of Seq2Seq models to the power prediction task mu2023improved . This means we can feed the neural network with time series of arbitrary length and prompt it to give us the forecast for any number of hours ahead. In this problem setting, the longer the input time series is, the better the model prediction is. Similarly, the shorter the required output length, the easier it is for the model to generate the forecast.
The seq2Seq model is a type of encoder-decoder model. The encoder is given the entire input sequence, which it uses to generate a context vector. This vector is used as an input hidden state for the decoder, so it literally provides it with “context,” according to which the decoder will generate the forecast. This means that the hidden dimensions of the encoder and the decoder must match.
The decoder creates the output sequence step by step. It starts with only the most recent entry. Based on this entry and the context vector, the decoder generates the second entry and appends it to the existing one. The obtained two-entry sequence is once again fed into the decoder to generate the third entry. The cycle then repeats until the length of the generated sequence matches the length requested by the user.
@@
In the QDI layer, instead of attempting to use a qubit for each feature asel1 , we employed the data re-uploading technique schuld2021effect ; data_reuploading . Specifically, we work with four qubits and structure them into a lattice of depth $4$ (depicted as a blue big rectangle in Fig. 3(d)). Each of our $16$ input features leading to the quantum layer is intricately encoded within this lattice. The first four features are mapped onto the initial depth, followed by the subsequent features in blocks of four. Encoding these classical features into the quantum domain, we perform “angle embedding” using $R_{z}$ gates. This operation effectively translates the input vector into a quantum state that symbolizes the preceding classical layer’s data.
@@
Entangling variational layers, signified by purple squares, are interspersed between every encoding layer, ensuring optimal Fourier accessibility. Each variational layer has two components: rotations governed by trainable parameters, and sequential CNOT gates. The rotations are implemented by quantum gates that manipulate the encoded input in line with the variational parameters, while the CNOT gates entangle the qubits.
@@
Each large blue bounding box encompasses a variational layer (small purple squares). Prior to all the encoding layers, we introduce a single variational layer (small green squares) for enhanced model representation. Consequently, the total weight count in the quantum segment of our network is $20$. In the measurement phase, all the qubits execute a CNOT operation where the first qubit is the target. This ensures that when the first qubit is measured in the $Y$-basis, the measurement outcome depends on the all qubits involved in the circuit. In this way, the output of the quantum layer serves as the power value prediction for a specific hour.
@@
The input size of the encoder and decoder can differ, which is a substantial benefit. For instance, we can use all of the $5$ features to create a context vector, but request to generate the forecast for only $1$ feature. Exploiting this advantage, we will feed the Seq2Seq model with a window of all known features and only demand the power forecast. For the sake of simplicity, we will train both models with a fixed input and output length of $96$ hours and then vary the length in the testing stage.
@@
Figure 4: Results of training and testing the HQNN, MLP, HQLSTM and LSTM models. (a-b) Training and testing history of the models shown by the dotted and solid lines, respectively. The filled space shows the standard deviation of that models averaged over the different testing subsets. (c) Histogram showing MAE and RMSE loss on the test dataset averaged over the results of 5 models trained on different subsets of training and test data. (d) The histogram shows MSE loss on the test dataset for MLP, HQNN, LSTM and HQLSTM models for the reduced training dataset size. (e) The histogram shows the RMSE loss on the test dataset for the Seq2Seq and HQSeq2Seq models for the reduced training dataset size. (f-g) The train and test learning curves for the Seq2Seq and HQSeq2Seq models. (h) An example of the classical and hybrid Seq2Seq models inference on the testing data. The models get $124$ hours of data as an input (before the dashed line) and give a forecast of the “Power” up to $137$ hours ahead (after the dashed line). The solid black line represents the ground truth value of the “Power” feature. The MSE/MAE errors in this particular example are $0.0052/0.0349$ for Seq2Seq and $0.0040/0.0292$ for HQSeq2Seq.